OpenStudy (anonymous):

simplify 2√6 * √5 + 4√6 *√5-√6 6√24 12√30 6√11 - √6 6√30 - √6

OpenStudy (anonymous):

What do all the terms have in common?

OpenStudy (anonymous):

there all factors??

OpenStudy (anonymous):

I only see the \[\sqrt{6}\]

OpenStudy (anonymous):

in all three

OpenStudy (anonymous):

yea i do to so does that mean all 3 of those are wrong

OpenStudy (anonymous):

nope... let me look

OpenStudy (anonymous):

ok sorry im really bad at this

OpenStudy (anonymous):

They don't really simplify... my best ans is \[\sqrt{6}(6\sqrt{5}-1)\]. But they ignore the last term at first and multiply \[\sqrt{5}\sqrt{6}=\sqrt{5*6}\]

OpenStudy (anonymous):

so... \[\sqrt{5*6}(2+4)-\sqrt{6}\]

OpenStudy (anonymous):

so its D

OpenStudy (danjs):

apply the exponent rules for combining the multiplied roots, \[\sqrt{a}*\sqrt{b}=\sqrt{a*b}\] combine terms, the root(6) is still left

OpenStudy (danjs):

\[2\sqrt{30}+4\sqrt{30}-\sqrt{6}\] \[6\sqrt{30}-\sqrt{6}\]

OpenStudy (danjs):

you can combine the same roots, not different ones, when they are added together

OpenStudy (anonymous):

yea i figured it out thanks you are the one that actually makes me understand this problem

OpenStudy (danjs):

just have to remember like 4 or 5 of them exponent rules , combining things right

OpenStudy (danjs):

becuase, recall a square root is the same as 'to the 1/2' power

OpenStudy (anonymous):

will do thanks